1. Researchers study the impact of internalized bias on patient care. Nurses complete a survey which determines a score for their internalized bias from 0-100. Without knowing the internalized bias score, patients are asked to rank the subsequent care they receive from the nurse as poor, fair, good or exemplary. In this study your independent variable is measured at what level of measurement? Ratio

Ordinal

Nominal

Interval

2. A study examines the relationship between interpersonal violence and health in women. They include an independent variable of exposure to violence from an intimate partner, other personal violence or both. The independent variable is measured at what level? Nominal

Ordinal

Interval

Ratio

3.Researchers study the impact of internalized bias on patient care. Nurses complete a survey which determines a score for their internalized bias from 0-100. Without knowing the internalized bias score, patients are asked to rank the subsequent care they receive from the nurse as poor, fair, good or exemplary. In this study your dependent variable is measured at what level of measurement? Nominal

Ordinal

Interval

Ratio

4. Researchers examined the relationship between interpersonal violence and health in college age women. They include an independent variable of exposure to violence from an intimate partner, other personal violence or both. The researchers wish to present a table showing the frequencies for each of these exposures. What might you suggest to best display the data?

Bar chart

Line graph

Grouped frequency chart

Scatterplot

5. Researchers examine the impact of Omega-3 fatty acids on vasomotor symptoms in menopausal women. Subjects in the study are given either a placebo (0g daily), 1g daily or 1.8g daily and asked to record the level of the vasomotor symptoms they experience on a scale of 0-10 each day. The level of vasomotor symptoms is an example of what type of data?

Ratio

Interval

Nominal

Ordinal

6. Patients on your unit complete a functional mobility test. The test determines how many seconds it takes to walk 20 feet unassisted. Your patient’s speed is reported as a Z score of 2.3. You would expect your patient to walk 20 feet unassisted:

A. Faster than the average patient on the unit

B. Slower than the average patient on the unit

C. 2.3 times faster than the slowest patient on the unit

D. Unable to determine

7. Your unit completes the internal controls module at your workplace. Two of the nurses on your unit were internal control specialists in their last place of employment and scored a perfect score on the exam while the rest of the nurses had an average score of 78 with a range of 72-84 on the test. When you look at the grade distribution you are not surprised to see:

A positive skew

A normal distribution

A negative skew

The mean, median and mode are equal

8. You have collected data about the fasting blood glucose (FBG) level of participants in your study. You are delighted to find that the variable is normally distributed. If the mean FBG is 82 and the standard deviation is 2.8 in what range would you expect to find the FBG of 68% of your study participants?

79.2-84.8

76.4 - 87.6

73.6 - 90.4

Unable to determine with the given information

9. Researchers examined the relationship between interpersonal violence and health in college age women. The researchers administered a psychological distress questionnaire which produced a score on a scale of 0-12. What would be an appropriate measure of central tendency to consider for this variable?

A. Any of these measures of central tendency would be appropriate

B. Mean

C. Median

D. Mode
Answer #1

1. Independent variable: Internalized bias on patient care which measures from 0 to 100

0 included in the score so it is Ratio level of measurement.

2. Independent variable: Interpersonal violence

It has levels so the level of measurement is Nominal

3. Dependent variable: Patient care.

It measures as poor, fair, good or exemplary, so this is Ordinal level of measurement.

4. Frequencies corresponding to each exposure are given, so the Bar chart is the best display of the data.

The bar chart represents the exposures on x-axis and frequencies on the y-axis.

5.

The level of vasomotor symptoms is measured on scale 1 to 10 so it is an ordinal scale of measurement.

6. Patient’s speed is reported as a Z score of 2.3

Z-score is positive that is the time is more than the expected time(meantime) that means the patient is slower than the average patient.

Option B is correct.

B. Slower than the average patient on the unit

7. The average score is 78 with a range of 72 - 84. The scores follow the normal distribution since some of the scores are above the average and some of the scores are below the average.

When you look at the grade distribution you are not surprised to see A normal distribution.

8. The variable fasting blood glucose is normally distributed with mean 82 and a standard deviation of 2.8

For normal distribution (that is bell-shaped) according to the Empirical rule, 68% of data falls within one standard deviation from mean that is between μ – 1 * σ αnd μ+1 και σ

u-1*0= 82 - 2.8 = 79.2

and u+1*0 = 82 + 2.8 = 84.8

The range for the 68% is 79.2 - 84.8

9. The scores measured on a scale of 0-12.

Which score is most repeated can check, so the mode is the best measure of central tendency for this variable.

Mean and median can not find for such data set.